OpenStudy (anonymous):

g(x)=___4___ 4-5x what is the correct domain?xdoes not equal 4/5,x does not equal zero,is a real number and x≥4/5 or x does not equal 4?

jimthompson5910 (jim_thompson5910):

the denominator 4-5x can't be zero if it were, then 4-5x=0 -5x=-4 x=-4/(-5) x=4/5 So if x=4/5, then the denominator is zero (which results in a division by zero error) So the domain is the set of all real numbers but x cannot equal 4/5

OpenStudy (anonymous):

awesome answer!Your a true math genious!